OpenStudy (anonymous):

find all points having a y-coordinate of -3 whose distance from the point (1,2) is 13

OpenStudy (whpalmer4):

Distance formula for distance between two points \((x_1,y_1),(x_2,y_2)\) is \[d = \sqrt{(x_2-x_1)^2+(y_2-y_1)^2}\] You know \((x_1,y_1)=(1,2)\) and \(y_2 = -3\) and \(d = 13\) Plug everything in and solve for \(x_2\)

OpenStudy (whpalmer4):

don't forget to take both positive and negative values of the square root!

OpenStudy (anonymous):

thanks , i know this but my last step is x^2-2x-143 , i can't factorise it cause of 143

OpenStudy (whpalmer4):

Complete the square, or use the quadratic formula?

OpenStudy (whpalmer4):

Or remember that 143 = 11*13

OpenStudy (anonymous):

oh okay thanks so much , i go try it now =)